# From High-SNR Radar Signal to ECG: A Transfer Learning Model with Cardio-Focusing Algorithm for Scenarios with Limited Data

### v1

@@ IV-A1 Dataset Collection and Preparation
The dataset contains a total of $80$-minute synchronous radar-ECG pairs collected for $5$ healthy subjects ($3$ men, $2$ women) in $2$ indoor scenarios as shown in Figure 6. The subjects are asked to sit causally and are allowed to change postures during data collection, and each data trial lasts for $1$ minute. The distance between radar and human body varies from $0.5-1.2$m, and a longer distance causes the decrease of signal SNR with a smaller portion of the space points containing useful cardiac features.
TI-AWR 1843 radar with $2$ Tx and $4$ Rx is used for data collection with $8$ virtual antenna channels created [24], and the radar configurations are listed in Table I with the name provided in TI mmWave-Studio interface. The signal will be sampled as $200$Hz, and only a band-pass filter from $0.5$ to $50$Hz and a differentiator are used for removing respiration noise because the radar signal extracted from CF points already has high SNR. Lastly, the ECG ground truth is collected using TI ADS1292, and the related ECG processing (e.g., smoothing and peak finding) is realized by NeuroKit2 python package [32].
Figure 6: Indoor scenarios for data collection.
@@
Start Frequency $77$GHz     Frequency Slope $65$MHz/$\mu$s
Idle Time $10\mu$s     Tx Start Time $1\mu$s
ADC Start Time $6\mu$s     ADC Samples $256$
Sample Rate $5000$kbps     Ramp End Time $60\mu$s
Start/End Chirp Tx $0/2$     No. of Chirp Loops $2$
No. of Frames $12000$     Frame Periodicity $5$ms
@@ IV-B1 Parameters for CFT Algorithm
The constraint $\Omega$ for the CF point search is centered at the initial state $E_{0}$ with a range of $0.4\times 0.2\times 0.4$m as illustrated in Figure 2(b). In addition, the initial grid and search region size should be adjusted to fit the real-life physical unit as $\Gamma_{k}=\gamma_{k}=0.1$m, and the size will be limited as $\Gamma_{k}\geq\gamma_{k}\geq 0.001$m to prevent an exhaustive search within a meaningless small space. At last, SNRd is set to $0.01$ for the desired MSE between normalized synthetic template and signal envelope, and $k_{max}$ is set to $100$.
@@ IV-B2 Deep Learning Model Training
The deep learning model adopts the same backbone, ECG decoder and hyperparameters as in our previous open-sourced work [8] coded in PyTorch and trained on NVIDIA RTX 4090 (24GB). The total training epoch is set to $100$ with batch size $8$, and a $5$-fold cross-validation training strategy is adopted to split the dataset to make the most of the limited dataset while excluding the testing data from the training phase. The parameter for balancing the weight of sparse penalty is set as $\lambda_{s}=0.1$.
@@ IV-C Methods for Comparison
The comparison is performed with the representative methods based on accumulation and clustering to extract high-SNR radar signal:
  • De-ViMo [3] is proposed for heart rate monitoring and is based on the accumulation of signals from various dimensions (e.g., chirps, antennas, spatial points) to enhance cardiac features while mitigating noise. In addition, De-ViMo also improves the rough localization by identifying the peaks in micro-motion frequency bands instead of the entire FMCW bands.
  • MMECG [9] requires the calculation of numerous points in 3D space and applies clustering algorithm to improve SNR. Then, a pattern-matching process is performed to learn the common pattern from the clustered result and select the best radar signal(s).
In addition, to the best of our knowledge, this is the first paper that leverages SSL with an appropriate pre-text task for pre-training of radar-based ECG recovery, and only the performance analysis will be provided without comparison with other methods.
@@ V-A1 Effectiveness of CFT Algorithm
In contrast, only a few range bins will contain useful cardiac features if the rough body location is far from CF points, especially when increasing the monitoring range. Therefore, the signal accumulation may enhance the noises as shown in Figure LABEL:sub@fig:vimo_bad while the signal clustering may also encounter a failure due to the lack of homogeneous cardiac signals as shown in Figure LABEL:sub@fig:MMECG_bad. However, The proposed CFT could precisely locate the CF point with good SNR subject to the designed signal template and DFO searching strategy, and the extracted radar signal still shows clear peaks as shown in Figure LABEL:sub@fig:bf_bad.
@@ V-A2 Impact of Monitoring Range
To evaluate the performance of different methods when increasing the monitoring range, the quality of extracted radar signals for all trials is evaluated in terms of:
  • Absolute Peak error between ECG R peaks and the dominant peaks for the first vibrations in radar signal.
  • Missed detection rate (MDR) to count the cardiac cycles with no peak detected or with the absolute peak error larger than $150$ms [9].
@@
Figure LABEL:sub@fig:pk_err_point and LABEL:sub@fig:pk_mdr_point illustrate the peak error and MDR for all $80$ trials with corresponding fitting curves indicating the mean peak error or MDR. All the methods show similar performance in the short range and experience certain degradation with respect to the increasing range. In particular, MMECG shows larger degradation and variance for longer-range cases because the rough localization based on FMCW signal processing cannot provide accurate cardiac location, and the resultant evaluated points may not capture useful information for clustering. In contrast, De-ViMo could provide a better cardiac location, and the accumulated results show better accuracy than MMECG, while the long-range monitoring still affects the quality because the accumulation is not robust to non-gaussian noise. At last, the proposed CFT could precisely focus on the CF point and keep tracking the high-SNR points during data collection, providing the best results with small variance for both peak error and MDR.
@@
Figure 8: Illustration of performance in terms of peak error and MDR: (a) - (b) Scattered points with fitting curves along range axis; (c) - (d) CDF plots for all trails.
@@
In addition, the cumulative distribution function (CDF) plots for all trials are shown in Figure LABEL:sub@fig:pk_err_cdf and LABEL:sub@fig:pk_mdr_cdf. The proposed CFT algorithm achieves the best peak error with a median value of $0.022$ sec, while DE-ViMo and MMECG have worse performances with larger median values of $0.028$ sec and $0.033$ sec, respectively. Similarly, the precise localization and tracking of CF point also reduces the MDR for CFT results with a median value of $14\%$, while DE-ViMo and MMECG may affected by the accumulated noise or inaccurate cardiac localization with the median MDR of $17\%$ and $23\%$, respectively.
TABLE II: Performance of supervised ECG recovery
Methods MSE (×$\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$
@@
CFT $\mathbf{0.82}$ $\mathbf{85.47\%}$ $\mathbf{7.61}$ $\mathbf{6.85\%}$
V-A3 Impact of Signal Quality on ECG Recovery
The signals extracted using different methods are used for supervised training to verify the impact of different input qualities on the ECG recovery task. The quality of the recovered ECG signal is assessed in terms of:
  • The morphological accuracy is measured using MSE and Pearson-correlation coefficient (PCC), with MSE sensitive to the peak deviation and PCC focusing on the similarity between the ECG patterns.
  • The accurate recovery of ECG R peaks is crucial to coarse cardiac features calculation (e.g., heart rate variability) and is measured by absolute R peak error and MDR.
Table II shows the performance of the deep learning model trained with datasets yielded by different methods. The training based on CFT dataset achieves the best results on both morphological accuracy (MSE$=0.0082$ and PCC$=85.47\%$) and R peak recovery (Peak Error$=7.61$ms and MDR$=6.85\%$), because the high-SNR inputs provide accurate peak locations with minor noise that affects the ECG pattern generation, as shown in Figure LABEL:sub@fig:radar_clean and LABEL:sub@fig:ecg_pred_clean.
In contrast, MMECG and De-ViMo cannot preserve the signal quality especially for long-distance cases, and the noisy inputs will prevent the deep learning mode from identifying the accurate position of ECG pieces, causing large peak error and MDR, as shown in Figure LABEL:sub@fig:radar_noise and LABEL:sub@fig:ecg_pred_noise. It is worth noticing that poor signal SNR causes more degradation in peak error than morphological accuracy, because the ECG patterns share a similar shape and can be learned from other cardiac cycles, while the peak recovery (detection) fully relies on the current radar input and can be ruined by noises.
@@ V-A3 Impact of Signal Quality on ECG Recovery
TABLE III: Performance of SSL with ablation study
Methods MSE (×$\times 10^{2}$) $\downarrow$ Sparsity $\downarrow$     MSE (×$\times 10^{2}$) $\downarrow$ Sparsity $\downarrow$
$\mathbf{100\%}$ Dataset     $\mathbf{80\%}$ Dataset
SSL w/o sp $0.91$ $0.36$     $0.96$ $0.41$
SSL with sp $0.82$ $0.20$     $0.85$ $0.22$
$\mathbf{60\%}$ Dataset     $\mathbf{40\%}$ Dataset
SSL w/o sp $1.43$ $0.44$     Failed
SSL with sp $0.92$ $0.26$     $0.98$ $0.31$
sp for sparse penalty
@@ V-B Performance of ECG Recovery using Transfer Learning
V-B1 Evaluations and Ablation Studies of SSL
@@
The SSR task is crucial in the proposed transfer learning framework to provide latent representation that assists the further ECG pattern recovery, reducing the demand for radar-ECG pairs in the fine-tuning stage. The results of SSL are shown in Table III in terms of MSE and sparsity to illustrate the former and latter part (without $\lambda_{s}$) in the loss function (10) for SSL training.
@@
The experiment is repeated for different dataset scales with the ablation study on the use of sparse penalty, and the results indicate that both MSE and sparsity decrease with the reducing training data as shown in Table III. Training with $100\%$ or $80\%$ dataset could achieve convergence and realize a successful SSR with similar MSE ($0.0091,0.0082$) or ($0.0096,0.0085$). However, the performance of SSR degrades heavily when further decreasing the training data without the constraint of sparse penalty, because the SSR results might fluctuate, as shown in Figure LABEL:sub@fig:ssr_bad. In contrast, introducing the sparse penalty could suppress the fluctuation and force the deep learning model to focus only on the dominant peaks of the input radar signals, as shown in Figure LABEL:sub@fig:ssr_good. Therefore, the training with sparse penalty loss still achieves good results with an MSE of $0.0092$ and $0.0098$ by using $60\%$ and $40\%$ of the dataset, while the model cannot be well-trained without sparse penalty by using $40\%$ dataset as shown in Table III and Figure LABEL:sub@fig:ssr_bad.
V-B2 Evaluations of Fine-tuning Results

### v2

@@ IV-A1 Dataset Collection and Preparation
The dataset contains a total of $80$-minute synchronous radar-ECG pairs collected for $5$ healthy subjects ($3$ men, $2$ women) in $2$ indoor scenarios as shown in Figure 6. The subjects are asked to sit casually and are allowed to change postures during data collection, and each data trial lasts for $1$ minute. The upper body travel distances vary from $1$ to 99m during 11-minute trials based on the statistics analyzed after the experiment. The distance between radar and human body varies from $0.5$ to 1.21.2m, and a longer distance causes a decrease in signal SNR, with a smaller portion of the space points containing useful cardiac features.
TI-AWR 1843 radar with $2$ Tx and $4$ Rx is used for data collection with $8$ virtual antenna channels created [24], and the radar configurations are listed in Table I with the names provided in TI mmWave-Studio interface. Meanwhile, the ECG ground truth is collected using TI ADS1292, with AC coupling and an integrated right-leg drive (RLD) amplifier to remove potential baseline drift or power-line noise. The related ECG processing (e.g., smoothing and peak finding) is realized by NeuroKit2 Python package [32]. All the signals are sampled at 200200Hz, and only a band-pass filter from $0.5$ to 5050Hz and a differentiator are used for removing respiration noise because the radar signal extracted from CF points already has high SNR. Lastly, the CFT algorithm is run on the Apple M2 Pro CPU with 3232GB RAM.
Figure 6: Indoor scenarios for radar signal collection, with the synchronous ECG ground truth collected.
@@
[…]
Start/End Chirp Tx $0/2$     No. of Chirp Loops 22
No. of Frames $12000$     Frame Periodicity 55ms
Tx Gain ($G_{T}$) $10$ dBi     Rx Gain ($G_{R}$) $30$ dBi
Tx power ($P_{\mathrm{T}}$) $12$ dBm     Noise Figure (NF) $15$ dB
Wavelength ($\lambda$) $3.9$ mm     Bandwidth (B) $3.8$ GHz
@@ IV-B1 Parameters for CFT Algorithm
The constraint $\Omega$ for the CF point search is centered at the initial state $E_{0}$ with a range of $0.4\times 0.2\times 0.4$m as illustrated in Figure 2(b). In addition, the initial grid and search region size should be adjusted to fit the real-life physical unit as $\Gamma_{k}=\gamma_{k}=0.1$m, and the size will be limited as $\Gamma_{k}\geq\gamma_{k}\geq 0.001$m to prevent an exhaustive search within a meaningless small space. Then, SNRd is empirically set to $0.01$ for the desired MSE between the normalized synthetic template and the signal envelope, ensuring that the extracted signal contains enough cardiac features for ECG recovery. At last, $k_{max}$ is empirically set to $100$ to balance the quality and cost of the space search.
@@ IV-B2 Deep Learning Model Training
The deep learning model adopts the same backbone, ECG decoder and hyperparameters as in our previous open-sourced work [8] coded in PyTorch and trained on NVIDIA RTX 4090 (24GB), and the total training epoch is set to $100$ with batch size $8$. All the experiments are repeated 5 times, and the mean values are taken to ensure the results are statistically reliable, but the confidence interval is not provided because all the improvements show statistical significance.
The dataset is split based on a 5-fold cross-validation training strategy, with the trials from 1 fixed subject for testing and the other 4 subjects alternately selected for training or validation, ensuring to make use of all the possible trials while not involving the testing data in the training phase. During each training iteration, the 4-sec input radar segment is split from each 1-minute trial with a step length of 1 sec, generating a total of 4560 samples for training, evaluation and testing.
TABLE II: Comparison table for different methods
Methods (Improve SNR) Feature Vital Sign Real- time Allow posture change     Methods (SSL) Input Output Loss
De-ViMo [3] Accumulation HR ✗     MaeFE [33] Masked radar signal Original radar signal MSE
MMECG [9] Clustering ECG ✗     CL [34] Pos. and neg. samples Latent vectors NT-Xent [34]
CFT DFO ECG ✔     RFcardi Radar signal Sparse heartbeats MSE
For the original De-ViMo with only HR monitoring ability.
@@ IV-C Methods for Comparison
The comparison is performed with the representative methods based on accumulation and clustering to extract high-SNR radar signals:
  • De-ViMo [3] is proposed for real-time heart rate (HR) monitoring and is based on the accumulation of signals from various dimensions (i.e., $2$ chirps, $8$ virtual antennas, $10$ spatial points) to enhance cardiac features while mitigating noise. However, original De-ViMo only generates a sinusoidal wave, which cannot be used for ECG recovery. Therefore, an extra signal extraction stage is needed, but the real-time monitoring will be unachievable.
  • MMECG [9] requires the calculation of numerous points in 3D space and applies clustering algorithm to improve SNR. Then, a pattern-matching process is performed to learn the common pattern from the clustered result and select the best radar signal(s). In this study, the number of 3-D points used for clustering is $6\times 6\times 6=216$.
In addition, to the best of our knowledge, this is the first paper that leverages SSL with an appropriate pre-text task for pre-training of radar-based ECG recovery. Therefore, two widely applicable SSL methods for biological signals are used for comparison:
  • MaeFE [33] is designed based on the famous masked autoencoder (MAE), with masked reconstruction of signal patches as the pre-text task. To fit our task, each 4-sec input radar signal will be masked using faint Gaussian noises ($3$ dB) with a length of $1$ sec.
  • The method proposed in [34] is based on contrastive learning (CL) by constructing positive and negative pairs and maximizing similarity between the latent vectors generated for positive pairs while minimizing similarity for negatives. To fit our task, the positive pairs are built by the records taken from two different channels in the same 4‑sec window, while the negative pairs are formed by pairing a window with another window more than 20 seconds apart in time, ensuring a different physiological state driven by changing activity.
The core features of all the methods used in this paper have been summarized in Table II for clarity.
IV-D Evaluation Metrics
The performance of the proposed CFT-RFcardi framework will be separately evaluated in terms of the quality (SNR) of the collected radar signal and the recovered ECG signal using the following metrics:
  • For both radar collection and recovered ECG signals, the peak accuracy is crucial because it links to many important cardiac features (e.g., heart rate variability) for clinical diagnosis. Thus, absolute peak error between ECG R peaks and the dominant peaks for the first vibrations (i.e., heartbeats) in the radar or recovered ECG signal is used to measure the peak accuracy.
  • Another metric related to peaks is missed detection rate (MDR) to count the cardiac cycles with no peak detected or with the absolute peak error larger than 150150ms [9], providing a quantified evaluation of the corrupted radar signal collection or ECG recoveries due to the noise.
  • For ECG recovery, the morphological accuracy is measured using MSE and Pearson-correlation coefficient (PCC) as calculated in (11), with MSE sensitive to the peak deviation and PCC focusing on the similarity between the ECG patterns.
    $\mathrm{MSE}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(x_{i}-x_{i}^{\prime}\right)^{2}$ (11)
    $\operatorname{PCC}$ $=\frac{\sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)\left(x_{i}^{\prime}-\bar{x}^{\prime}\right)}{\sqrt{\sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}}\sqrt{\sum_{i=1}^{N}\left(x_{i}^{\prime}-\bar{x}^{\prime}\right)^{2}}}$
    where the predicted and ground truth ECG pieces are $x$, $x^{\prime}$, with length $N$ and mean values $\bar{x}$, $\bar{x}^{\prime}$.
@@ V-A1 Effectiveness of CFT Algorithm
In contrast, only a few range bins will contain useful cardiac features if the rough body location is far from CF points, especially when increasing the monitoring range. Therefore, the signal accumulation may enhance the noises as shown in Figure LABEL:sub@fig:vimo_bad while the signal clustering may also encounter a failure due to the lack of homogeneous cardiac signals as shown in Figure LABEL:sub@fig:MMECG_bad. However, The proposed CFT can precisely locate the CF point with good SNR subject to the designed signal template and DFO searching strategy, and the extracted radar signal still shows clear peaks as shown in Figure LABEL:sub@fig:bf_bad.
@@ V-A2 Impact of Monitoring Range
To evaluate the performance of different methods when increasing the monitoring range, the quality of extracted radar signals for all trials is evaluated in terms of peak error and MDR. Figure LABEL:sub@fig:pk_err_point and LABEL:sub@fig:pk_mdr_point illustrate the peak error and MDR for all $80$ trials, with fitting lines indicating the mean peak error or MDR.
All the methods show similar performance in the short range and experience certain degradation with respect to the increasing range, because the SNR naturally decreases as the monitoring range increases. Following the link budget analysis for a similar radar platform used in [7], the degradation in SNR is around $14$ dB when the monitoring range increasing from 0.50.5m to 1.11.1m, causing a similar trend of degradation as shown in Figure LABEL:sub@fig:pk_err_point and LABEL:sub@fig:pk_mdr_point.
@@
In particular, MMECG shows larger degradation and variance for longer-range cases because the rough localization based on FMCW signal processing cannot provide accurate cardiac location, and the resultant evaluated points may not capture useful information for clustering. In contrast, De-ViMo could provide a better cardiac location, and the accumulated results show better accuracy than MMECG, while the long-range monitoring still affects the quality because the accumulation is not robust to non-Gaussian noise. At last, the proposed CFT could precisely focus on the CF point regardless of the monitoring range, providing the best results with small variance for both peak error and MDR.
@@
Figure 8: Illustration of performance in terms of peak error and MDR from different perspectives: (a) - (b) Scattered points and fitting lines w.r.t monitoring range; (c) - (d) Scattered points and fitting lines w.r.t moving distance of the CF points; (e) - (f) CDF plots for all trials.
V-A3 Impact of Posture Change
As one of the core innovations, CFT could locate and keep tracking the CF point to ensure a high-quality signal acquisition after the posture change. To provide a quantified evaluation, the posture change is counted as the travel distance of the CF points during each $1$-minute trial to measure the extent of upper body movement, and the results are shown in Figure LABEL:sub@fig:pk_err_point_dis and LABEL:sub@fig:pk_mdr_point_dis.
Similar to the impact of range, the peak error and MDR increase with moving distance because the signal collected during the posture change is heavily ruined, also known as random body movement distortion [9]. MMECG and De-ViMo will be affected if the CF points deviate too much from the initial body location and cannot guarantee a high SNR, because they do not have enough features for clustering or accumulation. In contrast, CFT could effectively find the new CF points after rounds of spatial search, and the extracted signal quality is better than the other two methods, as indicated by the fitting lines in Figure LABEL:sub@fig:pk_err_point_dis and LABEL:sub@fig:pk_mdr_point_dis.
V-A4 Signal Qualities for All Trials
@@
The cumulative distribution function (CDF) plots for all trials are shown in Figure LABEL:sub@fig:pk_err_cdf and LABEL:sub@fig:pk_mdr_cdf. The proposed CFT algorithm achieves the best peak error with a median value of $0.022$ sec, while DE-ViMo and MMECG have worse performances with larger median values of $0.028$ sec and $0.033$ sec, respectively. Similarly, the precise localization and tracking of CF point also reduces the MDR for CFT results with a median value of $14\%$, while DE-ViMo and MMECG may be affected by the accumulated noise or inaccurate cardiac localization with the median MDR of $17\%$ and $23\%$, respectively.
V-A5 Evaluations of Computational Cost
The outstanding performance of CFT owes to the space search mechanism that triggers new rounds of search if the current points no longer satisfy the desired SNR (SNRd). In comparison, MMECG and De-ViMo perform a fixed number of evaluations for every iteration, wasting a huge amount of time to calculate redundant points or channels. For all three methods, the most time-consuming step is to extract signals from points as calculated in (1) and (2), while the time consumed for clustering, accumulation, and space search is negligible. Therefore, the quantitative computational cost can be estimated by counting the evaluation times for each 4-sec signal segment:
  • MMECG takes $216$ spatial points for clustering, and the total number of evaluations for a $1$-minute trial is $3240$.
  • De-ViMo needs to consider the signals extracted from $2$ chirps, $8$ virtual antennas, $10$ spatial points, and the number of evaluations is $2400$.
  • CFT only evaluates the point during space search to get the cost values, and the number of evaluations recorded for each trial is shown in Figure 9.
According to the experiment, each evaluation takes about 6565ms, and the overall time consumed for evaluating a $1$-minute trial is also shown in Figure 9. Apparently, the proposed CFT saves a huge amount of time compared with other methods, and the time used for space search increases if the subjects change their posture more frequently (i.e., with longer moving distance), as shown in Figure LABEL:sub@fig:evaluation_num_rag. In addition, there is no obvious relationship shown between evaluation numbers and monitoring ranges as shown in Figure LABEL:sub@fig:evaluation_num, indicating that the CFT algorithm can keep extracting high-quality radar signals regardless of monitoring range and could save huge computational resources if the subject is static.
Figure 9: Number of evaluations and overall time consumed for each trial w.r.t the moving distance and monitoring range.
V-A6 Impact of Signal Quality on ECG Recovery
[…]
In contrast, MMECG and De-ViMo cannot preserve the signal quality especially for long-distance cases, and the noisy inputs will prevent the deep learning model from identifying the accurate position of ECG pieces, causing large peak error and MDR, as shown in Figure LABEL:sub@fig:radar_noise and LABEL:sub@fig:ecg_pred_noise. It is worth noticing that poor signal SNR causes more degradation in peak error than morphological accuracy, because the ECG patterns share a similar shape and can be learned from other cardiac cycles, while the peak recovery (detection) fully relies on the current radar input and can be ruined by noises.
TABLE III: Performance of supervised ECG recovery
Methods MSE ($\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$
@@ V-A6 Impact of Signal Quality on ECG Recovery
CFT $\mathbf{0.80}$ $\mathbf{85.47\%}$ $\mathbf{7.61}$ $\mathbf{6.85\%}$
@@ V-B Performance of ECG Recovery using Transfer Learning
V-B1 Evaluations and Ablation Studies of SSL in RFcardi
@@
The SSR task is crucial in the proposed transfer learning framework to provide latent representations that assist further ECG pattern recovery, reducing the demand for radar-ECG pairs in the fine-tuning stage. The results of SSL are shown in Table IV in terms of MSE and sparsity to illustrate the former (MSE term) and latter part (sparse penalty $\lambda_{s}$) in the loss function (10) for SSL training.
@@
The experiment is repeated for different dataset scales with the ablation study on the use of sparse penalty, and the results indicate that both MSE and sparsity decrease with the reduction in training data as shown in Table IV. Training with $100\%$ or $80\%$ dataset could achieve convergence and realize a successful SSR with similar MSE ($0.0091,0.0082,0.0081$) or ($0.0096,0.0085,0.0088$). However, the performance of SSR degrades heavily when further decreasing the training data without the constraint of sparse penalty ($\lambda_{s}=0$), because the SSR results might fluctuate, as shown in Figure LABEL:sub@fig:ssr_bad. In contrast, introducing the sparse penalty could suppress the fluctuation and force the deep learning model to focus only on the dominant peaks of the input radar signals, as shown in Figure LABEL:sub@fig:ssr_good. Therefore, the training with sparse penalty loss still achieves good results with an MSE of $0.0092$ and $0.0098$ by using $60\%$ and $40\%$ of the dataset, while the model cannot be well-trained without sparse penalty by using $40\%$ dataset as shown in Table IV and Figure LABEL:sub@fig:ssr_bad.
It is also revealed in Table IV about the consequence of enlarging the sparse penalty $\lambda_{s}$. Considering that even the ground truth shown in Figure 11 has a sparsity of $0.05$, using a large coefficient (e.g., $\lambda_{s}=0.05$) will overwhelm the MSE term, causing the DL model to generate the result with only $1$ non-zero value to give a sparsity of 0. In Table IV, $\lambda_{s}=0.02$ gives a higher MSE with lower sparsity compared with $\lambda_{s}=0.01$, especially when using $60\%$ and $40\%$ of the dataset (MSE=0.0118=0.0118 and $0.0124$, sparsity=0.024=0.024 and $0.026$). In other words, using a large sparse penalty cannot help the convergence when using a limited dataset, but may disrupt the training process due to the dominance of the sparse penalty term, causing suboptimal performance in capturing cardiac-related features (i.e., heartbeats).
TABLE IV: Performance of SSL and fine-tuning with ablation study using different $\lambda_{s}$
Methods SSL Fine-tuning Overall $\uparrow$
MSE ($\times 10^{2}$) $\downarrow$ Sparsity $\downarrow$ MSE ($\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$
$\mathbf{100\%}$ Unlabeled $\mathbf{100\%}$ Labeled
RFcardi with $\lambda_{s}=0$ 0.91{0.91} 0.36{0.36} $0.81$ $85.35\%$ $8.46$ $5.51\%$ $1.75\%^{1}$
RFcardi with $\lambda_{s}=0.01$ 0.82{0.82} 0.20{0.20} $0.80$ $85.51\%$ $8.40$ $5.14\%$ $\mathbf{3.66\%}$
RFcardi with $\lambda_{s}=0.02$ 0.81{0.81} 0.14{0.14} $0.80$ $86.60\%$ $8.34$ $5.47\%$ $2.97\%$}
$\mathbf{80\%}$ Unlabeled $\mathbf{80\%}$ Labeled
RFcardi with $\lambda_{s}=0$ $0.96$ $0.41$ $0.88$ $82.69\%$ $9.97$ $7.15\%$ $12.18\%$
RFcardi with $\lambda_{s}=0.01$ $0.85$ $0.22$ $0.82$ $84.53\%$ $9.95$ $6.94\%$ $\mathbf{-8.92\%}$
RFcardi with $\lambda_{s}=0.02$ $0.88$ $0.17$ $0.83$ $83.23\%$ $8.71$ $8.17\%$ $10.02\%$
$\mathbf{60\%}$ Unlabeled $\mathbf{60\%}$ Labeled
RFcardi with $\lambda_{s}=0$ $1.43$ $0.44$ $0.90$ $80.45\%$ $9.68$ $9.44\%$ $20.85\%$
RFcardi with $\lambda_{s}=0.01$ $0.92$ $0.26$ $0.91$ $81.56\%$ $9.31$ $7.58\%$ $\mathbf{-12.83\%}$
RFcardi with $\lambda_{s}=0.02$ $1.18$ $0.24$ $0.94$ $82.88\%$ $8.17$ $9.7\%$ $17.37\%$
$\mathbf{40\%}$ Unlabeled $\mathbf{40\%}$ Labeled
RFcardi with $\lambda_{s}=0$ - - - - - - Failed2
RFcardi with $\lambda_{s}=0.01$ $0.98$ $0.31$ $0.99$ $76.87\%$ $11.09$ $8.85\%$ $\mathbf{-27.18\%}$
RFcardi with $\lambda_{s}=0.02$ $1.24$ $0.26$ $1.13$ $66.43\%$ $12.23$ $10.41\%$ $44.05\%$
1The benchmark with $0\%$ improvement is given in Table V for supervised learning.
2The ECG recovery fails if $PCC<60\%$, according to the empirical observation of the morphological ECG features.
V-B2 Evaluations and Ablation Studies of Fine-tuning Results